OpenStudy (anonymous):

Please help!! Will medal and fan!!! The function for the cost of materials to make a hat is f(x) = one half x + 1, where x is the number of hats. The function for the selling price of those hats is g(f(x)), where g(x) = x + 2. Find the selling price of 10 hats. 6 8 10 12

OpenStudy (anonymous):

I do not understand how to set up an equation for g(f(x))

OpenStudy (lynfran):

g(f(x))... we input the value of the function f(x) where ever u see x in the g(x) function so g(f(x))=(one half x+1)+2

OpenStudy (anonymous):

I am not one of those people to ask the answer so, I wondering if you could give me an example so I can better understand.

OpenStudy (anonymous):

would the set up of the problem end up being: (1/2 times 10 + 1) +2 ?

OpenStudy (anonymous):

*sigh* help

OpenStudy (math&ing001):

Here's an example f(x) = 5x+8 and g(x) = 2x+1 g(f(x)) = g(5x+8) = 2(5x+8) +1 = 10x + 16 +1 = 10x + 17 Tell me what you get for your function.

OpenStudy (lynfran):

o my im now seeing this

OpenStudy (lynfran):

here another example f(x)=2x+5 g(x)=x-6 then g(f(x))= (2x+5)-6 and we solve this we get 2x-1

OpenStudy (lynfran):

so f(x)=1/2(x)+1 and we know x=10 so we can write f(10)= 1/2(10)+1 ==> 5+1==>6 so g(f(x))=(6)+2==>??